# Deep Unfolding of Fixed-Point based Algorithm for Weighted Sum Rate Maximization

### v1

@@ Abstract
In this paper, we propose a novel approach that harnesses the standard interference function, specifically tailored to address the unique challenges of non-convex optimization in wireless networks. We begin by establishing theoretical guarantees for our method under the assumption that the interference function exhibits log-concavity. Building on this foundation, we develop a Primal-Dual Algorithm (PDA) to approximate the solution to the Weighted Sum Rate (WSR) maximization problem. To further enhance computational efficiency, we leverage the deep unfolding technique, significantly reducing the complexity of the proposed algorithm. Through extensive numerical experiments, we demonstrate the competitiveness of our method compared to the state-of-the-art fractional programming benchmark, commonly referred to as FPLinQ.
@@ I Introduction
With the advent of Open Radio Access Network (OpenRAN) architectures for 6G, which aim to enable flexible, interoperable, and software-defined solutions, there is a pressing need for scalable and efficient algorithms that can address the complex interference management challenges in these emerging networks. In this paper, we propose a novel approach that integrates the standard interference function framework with difference-of-convex programming to address the challenging problem of maximizing the Weighted Sum Rate (WSR) in a Gaussian interference channel. This problem, which involves optimizing power control strategies, is inherently non-convex due to the complex interplay of interference in wireless networks. Our approach aims to overcome these challenges while providing scalable and efficient solutions suitable for next-generation wireless systems.
@@
Optimization algorithms can be unfolded into a finite number of iterations, treating these stages as layers of a deep neural network, with parameters trainable via stochastic gradient methods. This approach, known as deep unfolding [6], has recently gained attention in device-to-device (D2D) communications, as demonstrated by the algorithm-unfolding-based network proposed in [12].
@@ II-C System Model and Problem Formulation
For practical reasons in the great majority of works on D2D-networks it is assumed that interference is treated as noise, e.g. [17, 18, 19, 14, 26, 5, 29, 2, 10, 16, 11], which corresponds to the affine linear interference function
@@
where $\sigma_{i}$ is the noise variance.
@@ III-A Special Case
s.t. $0<{\bf p}_{i}\leq P_{\text{max}}\,\forall i=1,\dots,K$ (10b)
@@
$=\nabla_{\bf p}\left(\sum_{i}^{K}w_{i}\log\left(\mathcal{I}_{i}({\bf p}^{(k)})\right)\right)$ (11b)
@@
In the following we show that (12) converges to a fixed point by using the standard interference function framework by Yates [28]. Furthermore, it is shown that the right-hand side in (12) is a standard interference function itself for an arbitrary standard and feasible interference function $\mathcal{I}(\cdot)$, if $\mathcal{I}(\cdot)$ is $\log$-concave and the gradient w.r.t. to $\bf p$ is scale invariant as a function in $\bf q$.
@@
See appendix. [^1] The appendix can be found in the extended version.

### v2

@@ Abstract
In this paper, we propose a novel approach that harnesses the standard interference function, specifically tailored to address the unique challenges of non-convex optimization in wireless networks. We begin by establishing theoretical guarantees for our method under the assumption that the interference function exhibits log-concavity. Building on this foundation, we develop a Primal-Dual Algorithm (PDA) to approximate the solution to the Weighted Sum Rate (WSR) maximization problem. To further enhance computational efficiency, we leverage the deep unfolding technique, significantly reducing the complexity of the proposed algorithm. Through numerical experiments, we demonstrate the competitiveness of our method compared to the state-of-the-art fractional programming benchmark, commonly referred to as FPLinQ.
This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
With the advent of Open Radio Access Network (OpenRAN) architectures for 6G, which aim to enable flexible, interoperable, and software-defined solutions, there is a pressing need for scalable and efficient algorithms that can address the complex interference management challenges in these emerging networks. In this paper, we propose a novel approach that integrates the standard interference function framework with difference-of-convex programming to address the challenging problem of maximizing the Weighted Sum Rate (WSR) in a Gaussian interference channel. This problem, which involves optimizing power control strategies, is inherently non-convex due to the complex interplay of interference in wireless networks. Our approach aims to overcome these challenges while providing efficient solutions suitable for next-generation wireless systems.
@@
Optimization algorithms can be unfolded into a finite number of iterations, treating these stages as layers of a deep neural network, with parameters trainable via stochastic gradient methods. This approach, known as deep unfolding [6], has recently gained attention in wireless communications, as demonstrated by the algorithm-unfolding-based network proposed in [12].
@@ II-C System Model and Problem Formulation
For practical reasons in the great majority of works it is assumed that interference is treated as noise, e.g. [17, 18, 19, 14, 26, 5, 29, 2, 10, 16, 11], which corresponds to the affine linear interference function
@@
where $\sigma_{i}$ is the additive white Gaussian noise (AWGN) variance.
@@ III-A Special Case
s.t. $0<{\bf p}_{i}\leq P_{\text{max}}\,\forall i=1,\dots,K.$ (10b)
@@
$=\nabla_{\bf p}\left(\sum_{i}^{K}w_{i}\log\left(\mathcal{I}_{i}({\bf p}^{(k)})\right)\right).$ (11b)
@@
In the following we show that (12) converges to a fixed point by using the standard interference function framework by Yates [28]. Furthermore, it is shown that the right-hand side in (12) is a standard interference function for an arbitrary standard and feasible interference function $\mathcal{I}(\cdot)$, if $\mathcal{I}(\cdot)$ is $\log$-concave and the gradient w.r.t. to $\bf p$ is scale invariant as a function in $\bf q$.
@@
See appendix. ∎